# Partnering with Generative AI: Experimental Evaluation of Human-Led and Model-Led Interaction in Human-AI Co-Creation

### v1

@@ 4.1. Procedure
After providing informed consent, participants were randomly assigned to one of the five experimental conditions. In each condition, the participants were provided with an adapted version of the AUT task (Guilford, 1967). That is, participants had to come up with a creative way to repurpose existing features of cars, such as steering wheels, brakes, cameras, sensors, screens, lights, or speakers—for entirely new uses beyond the original purpose in vehicles. We chose this task to avoid the risk of public benchmarks of training data contamination, which can reflect model performance via memorization rather than generalization. Furthermore, it is an ecologically valid and complex task that is used exactly like this in companies. We therefore focused our task on combining novelty with usefulness in a highly realistic task.
@@ 4.2. Intervention Groups
Participants in both the model-led and human-led conditions first generated an initial idea without any LLM assistance. They then engaged with the chatbot according to their assigned condition. To ensure comparability, all LLM-assisted conditions used GPT-4.1 via the OpenAI API (snapshot gpt-4.1-2025-04-14). We followed best practices for the prompt design in the model-led and human-led condition (OpenAI, 2024). The full list of prompts is provided in Appendix A. In total, we implemented five conditions:
  • Model-led LLM ( M): Participants provided an initial idea (max. 10 words). A system prompt instructed the model to automatically improve the idea’s originality, appropriateness, and elaborateness without further user input.
@@ 4.3. Sample
Overall, N = 502 participants from the U.K. were recruited via the platform Prolific, as preregistered following an a priori sample size calculation. Inclusion criteria for our experiment were a minimum age of 18, English as a first language, an approval rate of at least 98%, and at least 50 previous submissions. The final sample consisted of participants with an age range of 18 to 78 years (M = 41.34, SD = 12.99); 49.21% identified as female, 50.61% as male, and one participant preferred not to disclose their gender. Following our preregistered exclusion criteria, we removed participants who failed at least one attention check (10 participants), acknowledged unauthorized AI use during the study (2 participants), provided nonsense answers (2 participants), participated more than once (1 participant), or did not finish the questionnaire (1 participant). This resulted in a final sample of N = 486 participants.
@@ 4.4. Measures
    Idea quality: To assess idea quality, we conducted the Consensual Assessment Technique (CAT), which is considered the gold standard to assess creativity (Amabile, 1982). As recommended, four domain experts—each with at least five months of experience in the innovation department of a major automotive manufacturer—rated all ideas independently on an originality and usefulness scale (the scale was adapted from Lee and Chung (2024)). For each idea, we averaged the originality and usefulness ratings across all raters to obtain a single idea quality score. The consistency between the raters reached an ICC (3,k) = .60 and is thus considered “good” according to Cicchetti (1994)
@@
    Diversity of the idea pool: We generated embeddings with OpenAI’s text-embedding-3-large model and, in line with Feuerriegel et al. (2025), measured per-idea diversity as the cosine distance to the leave-one-out centroid (mean of all other embeddings) within the same experimental condition.
@@
Beyond the primary outcomes, we examined additional variables as part of an exploratory analysis. To measure perceived cognitive workload, we adapted two items from the NASA-TLX (Hart, 2006), namely, perceived mental demand and effort. Participants provided their ratings on a 0–100 slider scale (e.g., “How mentally demanding was the task?”). The two items were strongly correlated ($r$ = .77). We further adopted three user experience-related scales, namely, performance expectancy, effort expectancy, and hedonic motivation (i.e., “I find the chatbot useful for creative tasks.”). These scales were adopted from the UTAUT questionnaire (Venkatesh et al., 2012) and internal consistency for the scales ranged from $\alpha$ = .92 to .97. We also assessed creative self-efficacy via three items adopted from Tierney and Farmer (2002) (i.e., “I was good at coming up with novel ways to repurpose vehicle features”; $\alpha$ = .89). Additionally, we performed an inductive qualitative content analysis (Hsieh and Shannon, 2005) to assess the participants’ elaboration depth in their interactions. Drawing on Guilford’s construct of elaboration (Guilford, 1967), as operationalized in the Torrance Tests of Creativity (Torrance, 1966), we treated elaboration as a sensitizing concept and transferred it to our interaction context. The first author open-coded the idea-refinement turns and, upon discussion with the other authors, iteratively organized incidents into seven characteristics for idea elaboration (CIEs) (i.e., adding a guiding principle, mechanism, feature, constraint, audience/context, integration, or reframing). Using these CIEs as decision rules, we derived a five-level ordinal scale of elaboration depth, ranging from Level 1 (AI-led echo, no new input) to Level 5 (human-led reframing, a radical pivot of the idea’s purpose or audience). This scheme allowed us to qualitatively assess the depth and originality of user contributions beyond simple acceptance of LLM output. The demographic measures were provided by the Prolific platform. For the context of our study, age and gender were included as demographic context variables in our analysis. Full measurement details, including all exploratory instruments and item wordings, are available in the Appendix.
@@ 4.5. Analysis
We analyzed each dependent variable using linear models with the condition as a predictor. Our preregistered one-sided hypotheses were tested via contrasts of estimated marginal means. To account for multiple comparisons across conditions, we adjusted $p$-values using the Holm method (Holm, 1979), thereby controlling the family-wise error rate. Statistical significance was set at $\alpha=.05$. Effect sizes are reported as Cohen’s $d$ (Cohen, 2013). In addition, each estimate is reported alongside its $95\%$ confidence interval (CI). All analyses were conducted in R 4.4.1, using the emmeans package (Lenth, 2025).
@@ 5.1.2. Idea Diversity (H2&H3): Significantly Higher for Question-Mode Condition
Figure 3. Distribution of idea diversity, idea quality, and perceived idea ownership across experimental conditions. Violin plots show the probability density of each measure, with wider sections indicating more frequent values. The boxplots show the interquartile ranges as well as the min/max, and the white dots indicate the medians.: Violin plots of idea diversity, idea quality, and idea ownership across conditions.
@@ 5.2. Influences of the Different Conditions on Perceived Cognitive Workload
Figure 4. Distribution of perceived cognitive workload across experimental conditions. Violin plots show the probability density of each measure, with wider sections indicating more frequent values. White dots indicate medians, thick black bars show interquartile ranges.: Violin plots of perceived cognitive workload across conditions.
In an exploratory analysis, we examined whether participants using the human-led LLMs perceived a higher cognitive workload than those using the model-led or vanilla LLM. To test this, we again conducted Holm-adjusted contrasts (detailed results are provided in Appendix G). Results showed that participants using the question-mode perceived significantly higher cognitive workload than participants using the model-led LLM ($t(481)=6.04$, $p<.001$, $d=0.86$) or the vanilla LLM ($t(481)=3.13$, $p=.004$, $d=0.45$). A similar pattern emerged for the suggestion-mode condition, which also yielded significantly higher workload compared to both model-led ($t(481)=5.47$, $p<.001$, $d=0.78$) and vanilla LLM ($t(481)=2.58$, $p=.010$, $d=0.37$). We further ran a linear regression predicting perceived ownership from cognitive workload, which was a significant positive predictor ($\beta=0.17$, $t(484)=3.83$, $p<.001$). However, the model explained only a small proportion of the variance in perceived ownership ($R^{2}=.03$).
@@ 5.3. Within-Subject Changes Regarding Idea Diversity
Figure 5. Within-subject analysis. (A) We compare the idea diversity of the initial ideas without interaction with the LLM with (B) the idea diversity of the refined ideas after interacting with the LLM: Illustration of a pre-post design.
@@
As an additional exploratory analysis, we examined within-subject changes in idea diversity across the human-led (question, suggestion) and model-led conditions (Figure 5). Idea diversity was calculated for each participant’s initial idea and compared with the diversity of their refined idea, both measured relative to the centroid of all initial ideas in that condition. [^1] Semantic similarity was calculated using cosine distance between OpenAI text-embedding-3-large vectors. Within-condition diversity measured each idea’s distance to the centroid of all other ideas within the same experimental condition (leave-one-out). For the refined ideas, we compared each refined idea’s embedding to the centroid of the initial ideas within the same condition. Using the same centroid for both stages, we evaluated how interaction with the different models changed idea diversity (detailed results in Appendix H). We used linear mixed-effects with planned contrasts of estimated marginal means to analyze changes in idea diversity. We found significant differences. Whereas for the model-led ($\Delta=-0.09$, 95% CI [0.11,0.07][-0.11,-0.07], t(295)=9.80t(295)=-9.80, $p<.001$) and suggestion-mode condition ($\Delta=-0.03$, 95% CI [0.04,0.01][-0.04,-0.01], t(295)=2.67t(295)=-2.67, $p=.006$), diversity decreased significantly, diversity significantly increased for the question-mode ($\Delta=0.07$, 95% CI $[0.05,0.09]$, $t(295)=7.20$, $p<.001$). In short, interaction with the question-mode expanded idea diversity, whereas both the model-led and suggestion-mode narrowed it.
@@
Figure 6. Within-subject in idea diversity from initial to refined ideas. Both initial and refined ideas are measured against the centroid of initial ideas within each condition to track within-subject movement (which we refer to as “diversity score”). Gray lines show individual participant trajectories, thick colored lines show condition averages, and the dotted line represents the overall average across all conditions.:
@@ 5.4. Analysis of the Interaction Mode Perceptions and Elaboration Depth
To obtain further insights into the user interactions, we qualitatively coded 20 conversations per condition to assess the elaboration depth of participants’ contributions. The majority of participants in the model-led condition tended to simply accept the model refinement without further elaboration (90% of participants). Interestingly, people in the suggestion-mode condition did not deeply elaborate on the LLM suggestions but rather added some details, like additional features, to their ideas, which were already suggested by the LLM (25% did not elaborate at all, 65% with minor elaboration). In the question-mode, however, people substantially elaborated based on the interaction with the LLM (55%), and 45% even went further and added an incrementally new idea. For instance, one participant initially proposed repurposing car cameras for video calls. After the LLM pointed out its low-light and infrared capabilities, the participant expanded the idea into using such cameras for threat detection and personal security. As depicted in Figure 7, the higher mental demand and effort observed in the human-led conditions translated into deeper idea elaboration only in the question-mode. Notably, this additional effort did not diminish user experience. Instead, in the question, we descriptively even observed higher hedonic motivation, along with greater effort and performance expectancy.
Figure 7. Comparative profile of user experience and cognitive demand. All measures standardized to a 1–7 scale; elaboration depth and NASA-TLX scores rescaled for comparability.:

### v2

@@ 4.1. Procedure
After providing informed consent, participants were randomly assigned to one of the five experimental conditions. In each condition, the participants were provided with an adapted version of the AUT task (Guilford, 1967). That is, participants had to come up with a creative way to repurpose existing features of cars, such as steering wheels, brakes, cameras, sensors, screens, lights, or speakers—for entirely new uses beyond the original purpose in vehicles. We chose this task to avoid the risk of public benchmarks of training data contamination, which can reflect model performance via memorization rather than generalization. Furthermore, it is an ecologically valid and complex task that is used exactly like this in companies. We therefore focused our task on combining novelty with usefulness in an innovation context.
@@ 4.2. Intervention Groups
Participants in both the model-led and human-led conditions first generated an initial idea without any LLM assistance. They then engaged with the chatbot according to their assigned condition. To ensure comparability, all LLM-assisted conditions used GPT-4.1 via the OpenAI API (snapshot gpt-4.1-2025-04-14). We followed best practices for the prompt design in the model-led and human-led condition (OpenAI, 2024). The full list of prompts is provided in Appendix  A. In total, we implemented five conditions. In all model-led and human-led conditions, the intervention aimed to (i) enhance the idea’s originality (exploration) and (ii) enhance its development (elaboration).
  • Model-led LLM ( M): Participants provided an initial idea (max. 10 words). By imposing a word limit, we ensured that the LLM-model performed the primary creative elaboration rather than merely refining an already developed idea. [^1] In the validation study, we implemented a different word limit (max. 15 words) that is the same for all conditions, including both human-led and model-led conditions. Still, we arrived at the same conclusions (see Section 6). A one-shot prompt was used to enhance originality and usefulness, after which participants could refine the model’s output before submitting.
@@
We validated the prompt manipulation through iterative testing. First, we conducted multiple short sessions with external testers to verify that each prompt condition consistently elicited the intended interaction patterns and that instructions were clearly understood. We iteratively refined prompts based on observed deviations from intended behavior. Second, we conducted a pilot study with 50 participants to validate the complete technical system, confirm that function calls were executed as designed, and ensure participants understood the different interaction modes. We analyzed all dialogues and adapted the prompts and code to address minor issues.
@@ 4.3. Sample
We approximated our required sample size using G*Power for a one-tailed two-sample $t$-test. To detect at least a medium effect size ($d$ = 0.5) with 80% power and $\alpha=0.0125$ (Bonferroni correction as a conservative proxy for Holm), we needed approximately 78 participants per group. To account for potential exclusions such as due to attention checks, we aimed for 100 participants per group as preregistered. Overall, N = 502 participants from the U.K. were recruited via the platform Prolific. Inclusion criteria for our experiment were a minimum age of 18, English as a first language, an approval rate of at least 98%, and at least 50 previous submissions. The final sample consisted of participants with an age range of 18 to 78 years (M = 41.34, SD = 12.99); 49.21% identified as female, 50.61% as male, and one participant preferred not to disclose their gender. Following our preregistered exclusion criteria, we removed participants who failed at least one attention check (10 participants), acknowledged unauthorized AI use during the study (2 participants), provided nonsense answers (2 participants), participated more than once (1 participant), or did not finish the questionnaire (1 participant). This resulted in a final sample of N = 486 participants.
@@ 4.4. Measures
    Idea quality: To assess idea quality, we conducted the Consensual Assessment Technique (CAT), which is considered the gold standard to assess creativity (Amabile, 1982). As recommended, four domain experts—each with at least five months of experience in the innovation department of a major automotive manufacturer—rated all ideas independently on an originality and usefulness scale (the scale was adapted from Lee and Chung (2024)). For each idea, we averaged the originality and usefulness ratings across all raters to obtain a single idea quality score. The consistency between the raters reached an ICC (3,k) = .60 and is thus considered “good” according to Cicchetti (1994).
@@
    Diversity of the idea pool: We generated embeddings with OpenAI’s text-embedding-3-large model and, in line with Feuerriegel et al. (2025), measured per-idea diversity as the cosine distance to the leave-one-out centroid (mean of all other embeddings) within the same experimental condition. This method has been found to align well with human judgments of conceptual diversity in previous studies (Anderson et al., 2024a; Beaty and Johnson, 2021; Dumas et al., 2021).
@@
Beyond the primary outcomes, we examined additional variables as part of an exploratory analysis. To measure perceived cognitive workload, we adapted two items from the NASA-TLX (Hart, 2006), namely, perceived mental demand and effort. Participants provided their ratings on a 0–100 slider scale (e.g., “How mentally demanding was the task?”). The two items were strongly correlated ($r$ = .77). We further adopted three user experience-related scales, namely, performance expectancy, effort expectancy, and hedonic motivation (i.e., “I find the chatbot useful for creative tasks.”). These scales were adopted from the UTAUT questionnaire (Venkatesh et al., 2012) and internal consistency for the scales ranged from $\alpha$ = .92 to .97. We also assessed creative self-efficacy via three items adopted from Tierney and Farmer (2002) (i.e., “I was good at coming up with novel ways to repurpose vehicle features”; $\alpha$ = .89). Additionally, we performed an inductive qualitative content analysis (Hsieh and Shannon, 2005) to assess the participants’ elaboration depth in their interactions. Drawing on Guilford’s construct of elaboration (Guilford, 1967), as operationalized in the Torrance Tests of Creativity (Torrance, 1966), we treated elaboration as a sensitizing concept and transferred it to our interaction context. The first author open-coded the idea-refinement turns and, upon discussion with the other authors, iteratively organized incidents into seven characteristics for idea elaboration (CIEs) (i.e., adding a guiding principle, mechanism, feature, constraint, audience/context, integration, or reframing). Using these CIEs as decision rules, we derived a five-level ordinal scale of elaboration depth, ranging from Level 1 (AI-led echo, no new input) to Level 5 (human-led reframing, a radical pivot of the idea’s purpose or audience). This scheme allowed us to qualitatively assess the depth and originality of user contributions beyond simple acceptance of LLM output. The detailed coding scheme can be found in the Appendix  E. The demographic measures were provided by the Prolific platform. For the context of our study, age and gender were included as demographic context variables in our analysis. Full measurement details, including all exploratory instruments and item wordings, are provided in the public GitHub repository.
@@ 4.5. Analysis
We analyzed each dependent variable using linear models with the condition as a predictor. Our preregistered one-sided hypotheses were tested via contrasts of estimated marginal means. To account for multiple comparisons across conditions, we adjusted $p$-values using the Holm method (Holm, 1979), thereby controlling the family-wise error rate. Statistical significance for all hypothesis tests was set at $\alpha=.05$. Effect sizes are reported as Cohen’s $d$ (Cohen, 2013). In addition, each estimate is reported alongside its $95\%$ confidence interval (CI). All analyses were conducted in R 4.4.1, using the emmeans package (Lenth, 2025) for calculating the contrasts. As dependent variables included subjective ratings and bounded measures (cosine similarity) we conducted non-parametric robustness checks for all dependent variables (see Appendix H).
@@ 5.1.2. Idea Diversity (H2&H3): Significantly Higher for Question-Mode Condition
Outcome distributions by condition for quality, diversity, and ownership Three side-by-side violin plots, one each for idea quality, idea diversity, and perceived idea ownership, shown across five conditions (question-mode, suggestion-mode, model-led, vanilla, control). Each plot includes an overlaid boxplot and median marker. Idea quality is highest for the model-led and question-mode conditions relative to the others. Idea diversity is lowest for the model-led condition and higher for question-mode and vanilla. Perceived idea ownership is highest for the control condition and next highest for question-mode, with lower values for the other chatbot conditions.
Figure 3. Distribution of idea diversity, idea quality, and perceived idea ownership across experimental conditions. Violin plots show the probability density of each measure, with wider sections indicating more frequent values. The boxplots show the interquartile ranges as well as the min/max, and the white dots indicate the medians.
Violin plots of idea diversity, idea quality, and idea ownership across conditions. The figure shows three side-by-side violin plots comparing outcomes across the five experimental conditions. The first plot displays idea diversity scores, where the vanilla LLM and the question-mode cluster around the highest median. The second plot shows idea quality, with the highest median for the question-mode and the model-led condition. The third plot shows perceived idea ownership, where the control condition and the question-mode have the highest median rating compared to the three other LLM-assisted conditions. Each violin illustrates the distribution shape, with white dots marking medians and thick black bars marking interquartile ranges.
@@ 5.2. Influences of the Different Conditions on Perceived Cognitive Workload
Perceived cognitive workload distributions across conditions Single violin plot of perceived cognitive workload on a zero to one hundred scale across five conditions (question-mode, suggestion-mode, model-led, vanilla, control), with boxplots and medians overlaid. Question-mode and suggestion-mode show higher workload distributions than model-led and vanilla. Model-led has the lowest median workload, while control and the human-led modes show higher medians and wider spreads.
Figure 4. Distribution of perceived cognitive workload across experimental conditions. Violin plots show the probability density of each measure, with wider sections indicating more frequent values. The boxplots show the interquartile ranges as well as the min/max, and the white dots indicate the medians.
Violin plots of perceived cognitive workload across conditions. The figure shows violin plots comparing perceived cognitive workload across the five experimental conditions. The human-led modes are higher in median workload than the control and vanilla conditions. Each violin depicts the distribution shape, white dots indicate medians, and thick black bars show interquartile ranges.
In an exploratory analysis, we examined whether participants using the human-led LLMs perceived a higher cognitive workload than those using the model-led or vanilla LLM. To test this, we again conducted Holm-adjusted contrasts (detailed results are provided in Appendix F). Results showed that participants using the question-mode perceived significantly higher cognitive workload than participants using the model-led LLM ($t(481)=6.04$, $p<.001$, $d=0.86$) or the vanilla LLM ($t(481)=3.13$, $p=.004$, $d=0.45$). A similar pattern emerged for the suggestion-mode condition, which also yielded significantly higher workload compared to both model-led ($t(481)=5.47$, $p<.001$, $d=0.78$) and vanilla LLM ($t(481)=2.58$, $p=.010$, $d=0.37$). Notably, these workload differences align with the intended interaction designs, suggesting that participants in human-led modes perceived higher cognitive workload compared to participants interacting with the model-led conditions. We further ran a linear regression predicting perceived ownership from cognitive workload, which was a significant positive predictor ($\beta=0.17$, $t(484)=3.83$, $p<.001$). However, the model explained only a small proportion of the variance in perceived ownership ($R^{2}=.03$).
@@ 5.3. Within-Subject Changes Regarding Idea Diversity
Conceptual within-subject comparison from initial to refined ideas Two-panel illustration of within-subject change. Panel A (pre) shows a participant writing an initial idea, with a thought bubble containing a cluster of points and one highlighted point representing the initial idea. Panel B (post) shows the participant after interacting with a chatbot and producing a refined idea, with the highlighted point shifted to a new position in the thought-bubble space to indicate change from initial to refined idea.
Illustration of a pre-post design. Panel A (Pre) shows a person writing down an initial idea, depicted as a single red dot among gray circles. Panel B (Post) shows the same person after interacting with an AI system, refining the idea. The refined idea is marked with a red icon, and the thought bubble illustrates the movement from the initial red dot to a more elaborated version, highlighting the cosine distance change of the ideas within the embedding space.
@@
Figure 5. Within-subject analysis. (A) We compare the idea diversity of the initial ideas without interaction with the LLM with (B) the idea diversity of the refined ideas after interacting with the LLM
@@
As an additional exploratory analysis, we examined within-subject changes in idea diversity across the human-led (question, suggestion) and model-led conditions (Figure 5). Idea diversity was calculated for each participant’s initial idea and compared with the diversity of their refined idea, both measured relative to the centroid of all initial ideas in that condition. [^2] Semantic similarity was calculated using cosine distance between OpenAI text-embedding-3-large vectors. Within-condition diversity measured each idea’s distance to the centroid of all other ideas within the same experimental condition (leave-one-out). For the refined ideas, we compared each refined idea’s embedding to the centroid of the initial ideas within the same condition. Using the same centroid for both stages, we evaluated how interaction with the different models changed idea diversity (detailed results in Appendix G). We used linear mixed-effects with planned contrasts of estimated marginal means to analyze changes in idea diversity. We found significant differences. Whereas for the model-led ($\Delta=-0.09$, 95% CI [0.11,0.07][-0.11,-0.07], t(295)=9.80t(295)=-9.80, $p<.001$) and suggestion-mode condition ($\Delta=-0.03$, 95% CI [0.04,0.01][-0.04,-0.01], t(295)=2.67t(295)=-2.67, $p=.006$), diversity decreased significantly, diversity significantly increased for the question-mode ($\Delta=0.07$, 95% CI $[0.05,0.09]$, $t(295)=7.20$, $p<.001$). In short, interaction with the question-mode expanded idea diversity, whereas both the model-led and suggestion-mode narrowed it.
Trajectories of diversity scores from initial to refined ideas Three small line plots, one per condition (question-mode, suggestion-mode, model-led). The horizontal axis shows idea stage (initial ideas to refined ideas) and the vertical axis shows diversity score. Thin gray lines show individual participant trajectories; a thick line shows the condition mean; a dotted horizontal line shows the overall mean across conditions. The mean diversity score increases from initial to refined ideas in question-mode, decreases in suggestion-mode, and decreases more strongly in model-led.
@@
Figure 6. Within-subject in idea diversity from initial to refined ideas. Both initial and refined ideas are measured against the centroid of initial ideas within each condition to track within-subject movement (which we refer to as “diversity score”). Gray lines show individual participant trajectories, thick colored lines show condition averages, and the dotted line represents the overall average across all conditions.
Line plot with three panels, one for each condition: Question-mode, Suggestion-mode, and Model-led. The y-axis shows diversity scores, and the x-axis shows stages from initial to refined ideas. Each panel includes many thin gray lines connecting initial and refined ideas for individual participants, and one thick colored line showing the average trend per condition. In Question-mode, the average line slopes slightly upward, showing an increase in diversity. In Suggestion-mode and Model-led, the average lines slope downward, showing a decrease in diversity. A horizontal dashed line marks the overall mean across conditions.
@@ 5.4. Analysis of the Interaction Mode Perceptions and Elaboration Depth
To obtain further insights into the user interactions, we qualitatively coded a randomly selected subset of 20 conversations per condition (80 ideas in total) to assess the elaboration depth of participants’ contributions. Two trained raters independently assessed all ideas. Initial inter-rater agreement was substantial (weighted $\kappa=0.628$). Disagreements were resolved through discussion to reach consensus, and these consensus scores were used in all subsequent analyses. The majority of participants in the model-led condition tended to simply accept the model refinement without further elaboration (90% of participants). Interestingly, people in the suggestion-mode condition did not deeply elaborate on the LLM suggestions but rather added some details, like additional features, to their ideas, which were already suggested by the LLM (25% did not elaborate at all, 55% with minor elaboration). In the question-mode, however, people substantially elaborated based on the interaction with the LLM (60%), and 40% even went further and added an incrementally new idea. For instance, one participant initially proposed repurposing car cameras for video calls. After the LLM pointed out its low-light and infrared capabilities, the participant expanded the idea into using such cameras for threat detection and personal security. As depicted in Figure 7, the question-mode successfully leveraged higher mental demand and effort to drive deeper idea elaboration. Notably, the additional cognitive workload that had to be spent on interacting with the question-mode did not diminish user experience. Instead, here we descriptively even observed higher hedonic motivation, along with greater effort and performance expectancy compared to the other conditions. In contrast, participants interacting with the suggestion-mode had high perceived cognitive workload that did not translate into high elaboration depth. To examine whether qualitative patterns linked to quantitative outcomes, we correlated elaboration depth with our dependent variables. Results showed that deeper elaboration was associated with greater idea diversity ($r=0.28$, $p=.032$), but not with quality or ownership (both $p>.15$). Hence, individuals who have a high elaboration depth do not experience the homogenizing effect of LLM interaction, resulting in diverse ideas.
Radar chart comparing user experience, workload, ownership, and elaboration across modes Radar chart with seven axes: performance expectancy, elaboration depth, perceived cognitive workload, creative self-efficacy, perceived ownership, hedonic motivation, and effort expectancy. Four polygons represent question-mode, suggestion-mode, model-led, and vanilla. Question-mode is highest on elaboration depth and perceived ownership and remains high on the user experience axes. Model-led shows the lowest elaboration depth and lower ownership than the human-led modes. Vanilla and suggestion-mode fall between these profiles across most dimensions.
Radar chart with four overlapping polygons, each representing an AI interaction condition: Question-mode, Suggestion-mode, Model-led, and Vanilla. Axes include Performance Expectancy, Elaboration Depth, Mental Effort, Mental Demand, Perceived Ownership, Hedonic Motivation, and Effort Expectancy. Question-mode shows higher scores across most axes, especially elaboration depth and ownership. Suggestion-mode and Vanilla have moderate levels across all dimensions, while Model-led scores lower on elaboration depth and perceived ownership.
Figure 7. Comparative profile of user experience and cognitive demand. The radar chart displays mean scores across seven dimensions for each interaction mode (distinguished by color). The user experience-related measures as well as perceived ownership and creative self-efficacy were measured on 7-point Likert scales. Elaboration depth (derived from coding levels 1–5) and perceived cognitive workload (mean of mental demand and mental effort, 0–100) were linearly rescaled to a 1–7 range to facilitate visual comparison.
6. Validation Study
6.1. Experiment Design
To test the robustness of our results and their generalizability to another domain, we conducted a preregistered replication study (https://aspredicted.org/de6md8.pdf). In this study, we focus on a different task that is less stylized and, further, more relevant to management practice. Specifically, we adapted the product ideation task from Girotra et al. (2023), asking participants to generate a product idea for a physical good that is targeted for university students in the United Kingdom, with a retail price of less than GBP 40. We selected this task to evaluate our interaction modes outside the typical AUT context and in a more realistic domain.
In this study, we restricted the initial idea to 15 words across all conditions (i.e., both model-led and human-led conditions) to improve comparability of conditions. We also implemented an additional model-led iterative condition. This condition mirrored the multi-step structure used in the suggestion-mode and question-mode: after submitting their initial idea, participants were provided three suggestions based on the initial idea in the exploration stage, then they moved on to the elaboration stage, where they could choose between the first choice and two further suggestions after being able to refine the final idea before submitting.
The dependent measures were identical to those used in the initial study. In this applied setting, we additionally assessed the potential future value creation of ideas. Following prior work, we added a single intention-to-buy item to the rating instrument (Girotra et al., 2023). We also included one item assessing perceived contribution (“self vs. AI”) using a 0–100 slider scale to capture how participants attributed authorship (see results in Appendix I). All adapted prompts used in this validation study are provided in Appendix A.
For the evaluation of idea quality and intention-to-buy, we recruited $N=320$ raters with active student status via Prolific (each rating 20 ideas). We used students as evaluators because they are representative of the target consumers and thus well suited to also judge the perceived value of the ideas (Kornish and Ulrich, 2014). Our final sample in the validation study consisted of $N=640$ participants across six conditions.
Table 4. Summary of hypothesis tests (Study 2). Reported are $z$-statistics for the ordinal mixed model (idea quality) and $t$-statistics with degrees of freedom for linear models (idea diversity and perceived ownership), along with associated Holm-adjusted $p$-values and Cohen’s $d$ effect sizes. indicates the hypothesis was supported ($p<0.05$); indicates it was not. Significance levels: $^{***}\!:\,p<0.001$; $^{**}\!:\,p<0.01$; $^{*}\!:\,p<0.05$.
Hyp. Comparison zz/$t$ (df) $p$ $d$ Supported?
Idea quality
H1a Q Question-mode >> C Control $5.59$ $001^{***}$ $0.49$
H1b S Suggestion-mode >> C Control $0.41$ .684.684 $0.04$
H1c M M-led (1-shot) >> C Control $6.38$ $001^{***}$ $0.50$
H1d M-I M-led (iter.) >> C Control $4.48$ $001^{***}$ $0.39$
H1e V Vanilla >> C Control 0.94-0.94 .826.826 0.08-0.08
Idea diversity
H2a Q Question-mode >> M M-led (1-shot) 3.01(634)3.01\;(634) $006^{**}$ $0.39$
H2b Q Question-mode >> M-I M-led (iter.) 1.01(634)1.01\;(634) .400.400 $0.14$
H2c Q Question-mode >> V Vanilla 3.35(634)-3.35\;(634) .999.999 0.48-0.48
H3a S Suggestion-mode >> M M-led (1-shot) 8.00(634)8.00\;(634) $001^{***}$ $1.04$
H3b S Suggestion-mode >> M-I M-led (iter.) 5.54(634)5.54\;(634) $001^{***}$ $0.79$
H3c S Suggestion-mode >> V Vanilla 1.11(634)1.11\;(634) .400.400 $0.16$
Perceived idea ownership
H4a Q Question-mode >> M M-led (1-shot) 3.76(634)3.76\;(634) $001^{***}$ $0.49$
H4b Q Question-mode >> M-I M-led (iter.) 9.32(634)9.32\;(634) $001^{***}$ $1.33$
H4c Q Question-mode >> V Vanilla 6.91(634)6.91\;(634) $^{***} 1.001.00$
H5a S Suggestion-mode >> M M-led (1-shot) 1.58(634)1.58\;(634) .058.058 $0.20$
H5b S Suggestion-mode >> M-I M-led (iter.) 7.36(634)7.36\;(634) $001^{***}$ $1.05$
H5c S Suggestion-mode >> V Vanilla 4.96(634)4.96\;(634) $001^{***}$ $0.71$
6.2. Analysis
For our validation study, we estimated linear mixed-effects models including random intercepts for both raters and ideas. This approach adjusts for systematic differences in rating severity across raters and isolates the variance attributable to the ideas themselves. As in the initial study, we evaluated our preregistered directional hypotheses at $\alpha=.05$, applied Holm-adjusted $p$-values to control the family-wise error rate, and used simple linear regressions for our analyses regarding idea diversity and perceived idea ownership.
6.3. Results
The validation study confirms the findings from our main study. As in the initial study, participants interacting with a question-mode (H1a: $z=5.59$, $p<.001$, $d=0.49$) and model-led (one-shot) interaction mode (H1c: $z=6.38$, $p<.001$, $d=0.50$) produced an idea with higher rated idea quality compared to participants without LLM support. Participants in the added model-led condition with an iterative approach also produced ideas with higher quality compared to participants in the control condition (H1d: $z=4.48$, $p<.001$, $d=0.39$), though the effect size was slightly smaller compared to the model-led one-shot condition.
For idea diversity, participants in the question-mode condition again produced ideas with higher diversity compared to participants in the model-led (one-shot) condition (H2a: $t(634)=3.01$, $p=.006$, $d=0.39$). In contrast to the initial experiment, participants in the suggestion-mode also produced ideas with higher diversity compared to participants in both model-led conditions (H3a: $t(634)=8.00$, $p<.001$, $d=1.04$ for one-shot; H3b: $t(634)=5.54$, $p<.001$, $d=0.79$ for iterative). Contrary to our expectations, participants interacting with the question-mode did not produce ideas with higher diversity compared to participants in the iterative model-led condition (H2b: $t(634)=1.01$, $p=.400$, $d=0.14$).
Regarding perceived idea ownership, participants interacting with both human-led modes showed significantly higher perceived ownership compared to those interacting with model-led and vanilla modes. Participants interacting with the question-mode LLM perceived higher ownership than those using both model-led one-shot (H4a: $t(634)=3.76$, $p<.001$, $d=0.49$) and iterative (H4b: $t(634)=9.32$, $p<.001$, $d=1.33$) approaches. Similarly, participants interacting with the suggestion-mode LLM showed higher ownership than those using the model-led iterative approach (H5b: $t(634)=7.36$, $p<.001$, $d=1.05$), though the difference with the model-led one-shot approach was marginally non-significant (H5a: $t(634)=1.58$, $p=.058$, $d=0.20$).
We further conducted an exploratory analysis examining purchase intent ratings to assess whether AI interaction modes influenced the perceived market viability of generated ideas (see Appendix I). Ideas created in interaction with the question-mode ($M=2.51$, $SD=0.58$) and model-led iterative condition ($M=2.50$, $SD=0.56$) showed the highest purchase intent ratings, while the suggestion-mode showed the lowest ($M=2.20$, $SD=0.53$), with model-led one-shot ($M=2.24$, $SD=0.49$), vanilla ($M=2.30$, $SD=0.55$), and control ($M=2.34$, $SD=0.62$) conditions falling in between.
In summary, the validation study successfully replicated all key findings from the main study in a different domain. Notably, in the validation study, participants interacting with both human-led modes perceived higher idea ownership compared to those using model-led approaches and participants interacting with the iterative model-led condition did not experience the same homogenization effect of ideas as in the one-shot model-led condition.